# article.bib

### v2

@@ 1.1 Research Background
In the rapidly evolving fields of neuroscience and artificial intelligence, the relationship between psychological phenomena and neural signals remains an unsolved mystery. Although our understanding of brain structure and neural activity has deepened, inferring individual psychological experiences from microscopic biological signals remains highly challenging. This involves both technical difficulties and philosophical paradoxes, such as whether psychological phenomena can be reduced to neural signals [luppi2021like]. This study proposes a new approach from EEG biofeedback to perceptual regulation, exploring the interaction between music rhythm and neural activity to understand and guide human psychological states.
@@
In contrast, this study aims to follow a different path. Drawing on Gestalt psychology theory, it treats perception as the core unit of psychological phenomena, breaking through traditional reductionist thinking. By introducing LOD theory, the study proposes a new deep learning method to dynamically adjust the processing levels of EEG signals, improving the accuracy of perceptual state recognition and providing new pathways for personalized psychological therapy. At the same time, innovative data collection and processing methods are adopted to advance the development of this field.
@@ 1.3 Inspiration and Reference from Gestalt Psychology
Gestalt psychology emphasizes the holistic nature of experience and behavior, asserting that the whole is greater than the sum of its parts. In visual perception, there are numerous organizational principles, such as the laws of proximity and similarity. In the field of deep learning for image recognition, this holistic perspective aligns to some extent with how neural networks process images by extracting features from the overall pixel distribution, similar to how the human visual system organizes visual information based on Gestalt principles [chen2024gestaltcomputationalmodelpersistent]. Additionally, the concept of "insight" in Gestalt psychology shares a similar underlying mechanism with the Transformer model’s ability to predict the next word in a sequence [10.3389/fcomp.2024.1419831].
@@ 2.1.1 Data Collection
This study focuses on collecting EEG data from the researchers themselves, aiming to explore individualized psychological perception and EEG characteristics from the perspective of Gestalt psychology. Given the uniqueness of individual psychological experiences, portable EEG devices were used for data collection to capture these unique EEG patterns and their underlying psychological structures.
The experiments were conducted in a quiet daily environment, with the researchers wearing TGAM chips for data collection. The TGAM chip plays a crucial role in this study’s data collection process. It not only records raw data in real time but also provides spectral information for $\alpha$, $\beta$, $\theta$, and $\delta$ waves, as well as two key indicators: relaxation and focus. These indicators range from 0 to 100, with higher values indicating greater relaxation or focus, providing an important basis for quantitatively assessing the psychological state of the subjects, see Table 1.
@@
Each data collection session lasted 6 to 8 minutes and included 200 rhythm loops. The rhythm speed could be adjusted between 90 and 140 BPM, with each adjustment made in increments of 5 BPM. For example, when the speed was set to 120 BPM, each rhythm loop lasted 2 seconds (60 seconds / 120 BPM = 0.5 seconds per beat, with 4 beats forming a loop, thus each loop lasting 2 seconds) [Yedukondalu2025]. Before each session, the researcher selected a seasonal label (spring, summer, autumn, winter) reflecting their current psychological state, which served as the training label for the deep learning model. A total of 418 independent data collection sessions were completed, accumulating a rich dataset of paired EEG and music rhythm data.
@@ 2.1.2 Data Preprocessing
First, the dataset was divided into training and test sets using an 80%-20% split. If necessary, a validation set could also be created. This division helps evaluate model performance and prevent overfitting. Each sample contains the following information:
  • EEG Data: Shape (N, 256, 4, 2), where N represents the number of samples.
  • Focus and Relaxation Scores: Shape (N, 2), recording the focus and relaxation scores for each sample.
@@
EEG data preprocessing is a critical step before model training. Raw EEG data was collected without frequency-domain transformation to preserve the maximum amount of information. The sampling frequency was set to 512 Hz, with 256 data points collected per rhythm loop, ensuring precise alignment with the temporal structure of the music rhythm.
@@
Since the TGAM dry electrode chip was used, EEG signals may be affected by noise and artifacts, such as eye movement artifacts (blinking) and power line interference (50/60 Hz). To ensure signal quality, low-pass filtering and Independent Component Analysis (ICA) techniques were employed to remove these artifacts, enhancing signal purity, which is crucial for improving model accuracy.
@@ 2.2 Software Design and Development
The software system incorporates several key elements. For audio materials and MIDI structure, MIDI drum materials generated by software such as BandInBox 2015, Addictive Drums 2, Studio Drummer, and EZdrummer 2 were used, totaling approximately 4GB. These materials cover a variety of music styles, including jazz, rock, and hip-hop, and exhibit a wide range of drum complexity, from simple to intricate. Each MIDI file is pre-labeled with important information such as BPM (beats per minute), drum density, and style tags. This enables the system to dynamically adjust music rhythms based on user feedback, providing a rich foundation for personalized music modulation based on user EEG data and psychological states.
The user interface and interaction design were implemented using Unity’s GUI system, featuring two adjustment axes: tension-relaxation and complexity-simplicity. Users can click buttons on the interface, such as "more tense," "more relaxed," "more complex," or "more simple," to adjust drum rhythms in real time. After each adjustment, the system automatically selects suitable drum combinations from the MIDI library and plays them using Unity’s audio engine, see Figure 2. This real-time interactive design not only enhances users’ ability to autonomously modulate music rhythms but, more importantly, creates diverse experimental conditions for collecting EEG data related to different psychological states through users’ active operations. This allows the research to more comprehensively explore the complex relationship between music rhythm and psychological states.
@@
Additionally, Unity’s built-in reinforcement learning module (e.g., ML-Agents) was ingeniously applied to optimize drum rhythm adjustment strategies. The model dynamically adjusts drum parameters based on user EEG data (including key indicators such as relaxation and focus) and subjective feedback (e.g., clicking the "adapt to current state" button). Through continuous learning and optimization, the system gradually guides users toward the target perceptual state, achieving intelligent interaction between the software system and users. This provides an innovative data-driven approach to music rhythm modulation for the research.
@@ 2.3 Data Collection and Label Assignment
To ensure the validity of the data and the high quality of model training, we carefully designed a dynamic label assignment mechanism that runs through the entire data collection process. This mechanism is of great significance for accurately reflecting changes in users’ perceptual states.
2.3.1 Initial State Selection
At the beginning of each experiment, the subject is required to select a seasonal label ("spring," "summer," "autumn," or "winter") that reflects their current psychological state as the initial perceptual state. Data collection then continues for 200 rhythm loops (approximately 6–8 minutes). During the data collection process, the drum rhythm speed can be flexibly adjusted between 90 and 140 BPM, with each adjustment made in increments of 5 BPM. This ensures that EEG data is collected under different rhythm conditions. For example, when the drum rhythm speed is set to 120 BPM, each rhythm loop lasts 2 seconds. This standardized collection process facilitates subsequent data processing and analysis.
2.3.2 Initial State Selection
The core part of this mechanism is the state transition and label assignment. If the subject remains in the initially selected perceptual state at the end of the experiment, all collected data will be explicitly labeled as 100% initial state. For example, if the initial state is "autumn" and no state change occurs by the end of the experiment, all data will be labeled as 100% autumn. However, if the subject transitions from the initial state (e.g., "autumn") to another state (e.g., "summer") during the experiment, the data labels will gradually change from 100% initial state to 100% target state, forming a linear gradient process [Craik_2019]. This dynamic label assignment method precisely records the trajectory of the subject’s perceptual state changes during the experiment, providing the subsequent deep learning model with rich, time-series-labeled information. This helps the model better learn and understand the patterns of perceptual state transitions [ding2024eeg].
@@ 2.3.3 Data Storage and Usage
Each experiment generates 50 data structures of size 256×4×2. Each data structure contains important information such as EEG features, audio features, relaxation and focus scores, and perceptual state labels. These data comprehensively record various physiological and psychological indicators of the subject during the experiment and accurately reflect changes in perceptual states through the labels. These rich data are used to train and validate the deep learning model, ensuring that the model can accurately identify and guide perceptual states. This provides a solid data foundation for the entire study.
@@ 2.4.1 Data Input and Preprocessing
  • The model input is a 256×4×2 tensor, representing the EEG and audio signals for each rhythm loop, where:
    • Time Dimension: 256 data points, ensuring alignment with the temporal structure of the music rhythm.
    • Loop Dimension: 4 rhythm loops form a complete cycle, capturing the periodic features of the rhythm.
    • Feature Dimension: 2, including EEG signals (Raw Data, $\alpha$, $\beta$, $\theta$, $\delta$ waves) and audio features (BPM, drum density).
  • Data Preprocessing Steps:
    • Normalization: Normalize EEG and audio features to unify data scales.
    • Smoothing: Use low-pass filtering to remove noise while retaining low-frequency EEG signals.
    • Data Augmentation: Introduce time-axis shifting and noise perturbation to enhance model robustness.
@@ 2.4.2 Feature Extraction and Fusion Based on LOD Theory
  • CNN Architecture: A multi-layer CNN is used to extract spatiotemporal features. The structure is as follows:
    • First Layer: 1×1 convolution to capture local features, reduce parameters, and output 32 channels with ReLU activation.
    • 3×3 convolution + max pooling to extract spatial features of time and audio signals, followed by ReLU activation and 2×2 max pooling.
    • 3×3 convolution + max pooling to further extract high-level features, capturing the complex relationship between EEG signals and music rhythms, followed by ReLU activation and 2×2 max pooling.
@@
    • Feature Fusion: After feature extraction by convolutional layers, EEG and audio features are concatenated along the channel dimension to form a unified feature representation.
    • LOD Application:
        Importance-Based LOD: Assign weights ($W_{i}$) based on feature importance to enhance the model’s sensitivity to key features. The formula is(1).
        Distance Threshold Method: Simplify signals through time-series pooling operations to simulate perceptual features at different time scales. The formula is(2).
        Dynamic LOD: Introduce reinforcement learning to dynamically adjust feature detail levels based on real-time conditions. The formula is(3).
@@ 2.4.3 Perceptual State Classification
  • Fully Connected Layer: Map the feature vectors extracted by the CNN to perceptual state classifications (spring, summer, autumn, winter). ReLU activation is used to introduce non-linearity.
  • Output Layer: Use the softmax function to generate probability distributions for each perceptual state. The output is the probability of each of the four states.
@@ 2.4.4 Model Training and Optimization
    where $L$ is the loss function, $N$ is the number of samples, $C$is the number of categories (4 in this study, corresponding to the four seasons), $y_{ij}$ is the true label of the $i$-th is the true label of the $p_{ij}$ is the model’s predicted probability of the $i$-th sample belonging to the $j$-th category.
@@
    • Divide the dataset into training and test sets using an 80%-20% split to prevent overfitting.
    • Use early stopping to halt training when the validation loss no longer decreases.
@@ 2.5 Reinforcement Learning Framework
The reinforcement learning module dynamically adjusts music rhythms to guide users from their current perceptual state to the target state. This study employs a Deep Q-Network (DQN) combined with features extracted by the CNN and user feedback to optimize rhythm adjustment strategies [10220592], see Figure 3.
@@ 2.5.1 State Space Definition
  • Focus and Relaxation Scores: Focus and relaxation scores provided by the TGAM chip, ranging from 0 to 100, indicating the user’s current psychological state.
@@
$s_{t}=[f_{\text{CNN}},\text{Attention}_{t},\text{Relaxation}_{t},\text{BPM}_{t},\text{Density}_{t},\text{Style}_{t}]$ (6)
where $fCNN$ represents the feature vector extracted by the CNN, $\text{Attention}_{t}$ and $\text{Relaxation}_{t}$represent the focus and relaxation scores, and $\text{BPM}_{t}$,$\text{Density}_{t}$ and $\text{Style}_{t}$ represent the current BPM, drum density, and style tags, respectively.

### v3

@@ 1.1 Research Background
In the rapidly evolving fields of neuroscience and artificial intelligence, the relationship between psychological phenomena and neural signals remains an unsolved mystery. Although our understanding of brain structure and neural activity has deepened, inferring individual psychological experiences from microscopic biological signals remains highly challenging. This involves both technical difficulties and philosophical paradoxes, such as whether psychological phenomena can be reduced to neural signals [luppi2021like]. This study proposes a new approach from electroencephalogram (EEG) biofeedback to perceptual regulation, exploring the interaction between music rhythm and neural activity to understand and guide human psychological states.
@@
In contrast, this study aims to follow a different path. Drawing on Gestalt psychology theory, it treats perception as the core unit of psychological phenomena, breaking through traditional reductionist thinking. By introducing Level of Detail (LOD) theory, the study proposes a new deep learning method to dynamically adjust the processing levels of EEG signals, improving the accuracy of perceptual state recognition and providing new pathways for personalized psychological therapy. At the same time, innovative data collection and processing methods are adopted to advance the development of this field.
@@ 1.3 Inspiration and Reference from Gestalt Psychology
Gestalt psychology emphasizes the holistic nature of experience and behavior, asserting that the whole is greater than the sum of its parts. In visual perception, there are numerous organizational principles, such as the laws of proximity and similarity. This holistic perspective is conceptually illustrated in Figure 1. In the field of deep learning for image recognition, this holistic perspective aligns to some extent with how neural networks process images by extracting features from the overall pixel distribution, similar to how the human visual system organizes visual information based on Gestalt principles [chen2024gestaltcomputationalmodelpersistent]. Additionally, the concept of "insight" in Gestalt psychology shares a similar underlying mechanism with the Transformer model’s ability to predict the next word in a sequence [10.3389/fcomp.2024.1419831].
@@ 1.5 Inspiration and Reference from LOD Theory
The remainder of this paper is organized as follows. Section Methods 2 details the experimental methodology, including data collection, preprocessing, software design, label assignment, and the proposed deep learning model architecture incorporating LOD principles, as well as the reinforcement learning framework for perceptual guidance. Section Results 3 presents the experimental results for both perceptual state classification and guidance tasks. Finally, Section Discussion 4 discusses the findings, experimental limitations, philosophical reflections inspired by the study, and future research directions.
@@ 2.1.1 Data Collection
This study adopts an exploratory approach, focusing on collecting EEG data from a single researcher acting as the subject. The aim is to explore individualized psychological perception and EEG characteristics from the perspective of Gestalt psychology, acknowledging the limitations of single-subject analysis regarding generalizability. Given the uniqueness of individual psychological experiences, a portable EEG device equipped with a ThinkGear ASIC Module (TGAM) chip was used for data collection to capture these unique EEG patterns and their underlying psychological structures, specifically from the FP1 electrode position.
The experiments were conducted in a quiet daily environment, with the researcher wearing the TGAM chip device for data collection. The TGAM chip plays a crucial role in this study’s data collection process. It not only records raw data in real time but also provides spectral information for $\alpha$, $\beta$, $\theta$, and $\delta$ waves, as well as two key indicators provided by the chip: Attention and Meditation. These indicators range from 0 to 100, with higher values indicating greater relaxation or focus, providing an important basis for quantitatively assessing the psychological state of the subject, see Table 1.
@@
Each data collection session lasted 6 to 8 minutes and included 200 rhythm loops. The rhythm speed could be adjusted between 90 and 140 BPM, with each adjustment made in increments of 5 BPM. For example, when the speed was set to 120 BPM, each rhythm loop lasted 2 seconds (60 seconds / 120 BPM = 0.5 seconds per beat, with 4 beats forming a loop, thus each loop lasting 2 seconds) [Yedukondalu2025]. Before each session, the researcher selected a seasonal label (spring, summer, autumn, winter) reflecting their current psychological state, which served as the training label for the deep learning model. A total of 418 independent data collection sessions were completed for this subject, accumulating a rich dataset of paired EEG and music rhythm data.
@@ 2.1.2 Data Preprocessing
First, the dataset, comprising data from the single subject, was divided into training and test sets using an 80%-20% split. If necessary, a validation set could also be created. This division helps evaluate model performance and prevent overfitting. Each sample contains the following information (where N denotes the total number of samples):
  • EEG Data: Shape (N, 256, 4, 2), representing the raw time-series EEG signal structures for each loop.
  • Attention and Meditation Scores: Shape (N, 2), recording the Attention and Meditation Scores scores provided by the TGAM chip for each sample.
@@
EEG data preprocessing is a critical step before model training. To preserve the maximum amount of temporal information, raw EEG data was collected without frequency-domain transformation. The sampling frequency was set to 512 Hz, with 256 data points collected per rhythm loop, ensuring precise alignment with the temporal structure of the music rhythm.
However, since the TGAM dry electrode chip was used, EEG signals could be affected by noise and artifacts, such as eye movement artifacts (blinking) and power line interference (50/60 Hz). To ensure signal quality and improve model accuracy, specific preprocessing steps were employed:
  • Low-pass filtering: A low-pass filter with a cutoff frequency of [Specify Cutoff Frequency, e.g., 45 Hz] was applied to remove high-frequency noise while preserving relevant EEG bands.
  • Artifact Removal: Independent Component Analysis (ICA) techniques were employed to identify and remove components associated with eye blinks and other artifacts, enhancing signal purity.
@@ 2.2 Software Design and Development
The software system incorporates several key elements. For audio materials and Musical Instrument Digital Interface (MIDI) structure, MIDI drum materials generated by software such as BandInBox 2015, Addictive Drums 2, Studio Drummer, and EZdrummer 2 were used, totaling approximately 4GB. These materials cover a variety of music styles, including jazz, rock, and hip-hop, and exhibit a wide range of drum complexity, from simple to intricate. Each MIDI file is pre-labeled with important information such as Beats Per Minute (BPM), drum density, and style tags. This enables the system to dynamically adjust music rhythms based on user feedback, providing a rich foundation for personalized music modulation based on user EEG data and psychological states.
The user interface and interaction design were implemented using Unity’s Graphical User Interface (GUI) system, featuring two adjustment axes: tension-relaxation and complexity-simplicity. Users can click buttons on the interface, such as "more tense," "more relaxed," "more complex," or "more simple," to adjust drum rhythms in real time. After each adjustment, the system automatically selects suitable drum combinations from the MIDI library and plays them using Unity’s audio engine, see Figure 2. This real-time interactive design not only enhances users’ ability to autonomously modulate music rhythms but, more importantly, creates diverse experimental conditions for collecting EEG data related to different psychological states through users’ active operations. This allows the research to more comprehensively explore the complex relationship between music rhythm and psychological states.
@@
Additionally, Unity’s built-in reinforcement learning (RL) module (specifically, the ML-Agents toolkit) was ingeniously applied to optimize drum rhythm adjustment strategies. The model dynamically adjusts drum parameters based on user EEG data (including key indicators such as Attention and Meditation scores) and subjective feedback (e.g., clicking the "adapt to current state" button). Through continuous learning and optimization, the system gradually guides users toward the target perceptual state, achieving intelligent interaction between the software system and users. This provides an innovative data-driven approach to music rhythm modulation for the research.
@@ 2.3 Data Collection and Label Assignment
To ensure the validity of the data and the high quality of model training, we carefully designed a dynamic label assignment mechanism that runs through the entire data collection process. This mechanism is of great significance for accurately reflecting changes in the user’s perceptual states.
2.3.1 Initial State Selection and Dynamic Labeling
At the beginning of each experiment, the subject (the researcher) was required to select a seasonal label ("spring," "summer," "autumn," or "winter") that reflected their current psychological state as the initial perceptual state. Data collection then continued for 200 rhythm loops (approximately 6-8 minutes). During the data collection process, the drum rhythm speed could be flexibly adjusted between 90 and 140 BPM, with each adjustment made in increments of 5 BPM. This ensured that EEG data was collected under different rhythm conditions. For example, when the drum rhythm speed was set to 120 BPM, each rhythm loop lasted 2 seconds. This standardized collection process facilitated subsequent data processing and analysis.
The core part of this mechanism is the state transition and label assignment during the session. If the subject remained in the initially selected perceptual state at the end of the experiment, all collected data would be explicitly labeled as 100% initial state. For example, if the initial state was "autumn" and no state change occurred by the end of the experiment, all data would be labeled as 100% autumn. However, if the subject transitioned from the initial state (e.g., "autumn") to another state (e.g., "summer") during the experiment, the data labels would gradually change from 100% initial state to 100% target state, forming a linear gradient process [Craik_2019]. This dynamic label assignment method precisely records the trajectory of the subject’s perceptual state changes during the experiment, providing the subsequent deep learning model with rich, time-series-labeled information. This helps the model better learn and understand the patterns of perceptual state transitions [ding2024eeg].
@@ 2.3.2 Data Storage and Usage
Each experiment generates 50 data structures of size 256×4×2. Each data structure contains important information such as EEG features, audio features, Attention and Meditation scores, and perceptual state labels. These data comprehensively record various physiological and psychological indicators of the subject during the experiment and accurately reflect changes in perceptual states through the labels. These rich data are used to train and validate the deep learning model, ensuring that the model can accurately identify and guide perceptual states. This provides a solid data foundation for the entire study.
@@ 2.4.1 Data Input and Preprocessing
  • The model input is a tensor with dimensions 256×4×2 for each perceptual segment processed, representing the EEG and audio signals associated with a cycle of rhythm loops. The dimensions signify:
    • Time Dimension: 256 data points per loop, ensuring alignment with the temporal structure of the music rhythm.
    • Loop Dimension: 4 consecutive rhythm loops forming a complete cycle, capturing periodic features.
    • Feature Dimension: 2 channels integrating the relevant EEG signal information and the corresponding audio features (e.g., rhythm characteristics like BPM and density) for that time point and loop. The exact composition of these integrated features is determined during the feature engineering process based on the raw data and derived metrics (like band power).
  • Data Preprocessing Steps performed before model training:
    • Normalization: Normalize EEG and audio features to unify data scales, preventing features with larger values from dominating the learning process.
    • Data Augmentation: Introduce time-axis shifting and noise perturbation to the training data. This enhances model robustness and generalizability by exposing it to a wider variety of signal variations.
@@ 2.4.2 Feature Extraction and Fusion Based on LOD Theory
  • CNN Architecture: A multi-layer CNN is used to extract spatiotemporal features. The structure involves:
    • First Layer: 1×1 convolution to capture local features, reduce parameters, and output 32 channels, followed by Rectified Linear Unit (ReLU) activation.
    • Second Layer: 3×3 convolution followed by max pooling to extract spatial features of time and audio signals. This is followed by ReLU activation and 2×2 max pooling to further extract high-level features, capturing the complex relationship between EEG signals and music rhythms.
@@
    • Feature Fusion: After feature extraction by the convolutional layers, the resulting EEG and audio features are concatenated along the channel dimension to form a unified feature representation before applying LOD mechanisms.
    • LOD Application: Different LOD concepts are applied to modulate the feature representation:
        Importance-Based LOD: Assign weights ($W_{i}$) based on learned feature importance to enhance the model’s sensitivity to key perceptual features. These weights are typically learned during the model training process, often guided by the task objective (e.g., classification accuracy) or specific attention mechanisms within the network. The operation is defined by $\hat{X}=X\cdot W_{i}$ (1), where $X$ represents the input features to this stage and $\hat{X}$ represents the weighted features.
        Distance Threshold Method (Simulated): Simplify signals through time-series pooling operations (e.g., max pooling or average pooling) to simulate perceptual features at different time scales, analogous to reducing detail based on distance. The formula is $\hat{X}=\text{Pool}(X,k)$ (2), where $X$ is the input data, $Pool$ represents the pooling operation, and $k$ is the pooling window size.
        Dynamic LOD: Introduce reinforcement learning or attention mechanisms to dynamically adjust feature detail levels based on real-time conditions, such as predicted task relevance or auxiliary inputs like Attention/Meditation scores ($E$). The formula is $\hat{X}=f(X)\cdot\sigma(W_{e}\cdot E)$ (3), where $f(X)$ represents the features extracted by the preceding layers (e.g., the CNN), $E$ represents auxiliary context like Attention/Meditation levels, $W_{e}$ is a learned weight matrix for the context, and $\sigma$ is an activation function.
@@ 2.4.3 Perceptual State Classification
  • Fully Connected Layer: Map the feature vectors extracted by the CNN and potentially modulated by LOD mechanisms to perceptual state classifications (spring, summer, autumn, winter). ReLU activation is used to introduce non-linearity.
  • Output Layer: Use the softmax function to generate probability distributions over the perceptual states. The softmax function ensures that the outputs are non-negative and sum to one, representing the model’s confidence for each of the four possible states. The final output is the probability of each state.
@@ 2.4.4 Model Training and Optimization
    where $L$ is the loss function, $N$ is the number of samples, $C$ is the number of categories (4 in this study, corresponding to the four seasons), $y_{ij}$ is a binary indicator (0 or 1) indicating if category $j$ is the correct classification for sample $i$, and $p_{ij}$ is the model’s predicted probability of the $i$-th sample belonging to the $j$-th category.
@@
    • Divide the dataset into training and test sets using an 80%- 20% split to prevent overfitting. A separate validation set (10-20
    • The model was trained for a maximum of 200 epochs with a batch size of 32.
    • Use early stopping to halt training when the validation loss no longer decreased for a predefined number of epochs (10 patience epochs), selecting the model weights that achieved the best validation performance.
@@ 2.5 Reinforcement Learning Framework
The reinforcement learning module dynamically adjusts music rhythms to guide the user from their current perceptual state to the target state. This study employs a Deep Q-Network (DQN) combined with features extracted by the CNN and user feedback to optimize rhythm adjustment strategies [10220592], see Figure 3.
@@ 2.5.1 State Space Definition
  • Attention and Meditation Scores: Attention and Meditation scores provided by the TGAM chip, ranging from 0 to 100, indicating the user’s current psychological state (Attention corresponding to focus, Meditation corresponding to relaxation).
@@
$s_{t}=[f_{\text{CNN}},\text{Attention}_{t},\text{Meditation}_{t},\text{BPM}_{t},\text{Density}_{t},\text{Style}_{t}]$ (6)
where $f_{\text{CNN}}$ represents the feature vector extracted by the CNN, $\text{Attention}_{t}$ and $\text{Meditation}_{t}$ represent the attention and meditation scores, and $\text{BPM}_{t}$, $\text{Density}_{t}$ and $\text{Style}_{t}$ represent the current BPM, drum density, and style tags, respectively.